OpenStudy (anonymous):

Solve for m: w = m - 5L

OpenStudy (anonymous):

Add +5L to both the sides, w + 5L = m so, m = w + 5L is the answer..

OpenStudy (anonymous):

m=w+5l w=m-5l +5l +5l ______________ w+5l=m

OpenStudy (anonymous):

thanks